OpenStudy (anonymous):

how do you solve this equation 2=√x/2

OpenStudy (anonymous):

Try squaring both sides.

OpenStudy (anonymous):

i did but i dont get it

OpenStudy (anonymous):

What did you get when you squared both sides? Also is the equation \[2 = \frac{\sqrt{x}}{2}\] or \[2 = \sqrt{(x/2)}\]

OpenStudy (anonymous):

its the 2nd one

OpenStudy (anonymous):

Ok, so squaring both sides you got?

OpenStudy (anonymous):

ok im going to do it again

OpenStudy (anonymous):

i got 4 but i dont know if that is right

OpenStudy (anonymous):

That is not correct. \[2 = \sqrt{(x/2)}\] \[\rightarrow 2^2 = \sqrt{(x/2)}^2\] \[\rightarrow 4 = \frac{x}{2}\] \[\rightarrow x = ?\]

OpenStudy (anonymous):

thank you i get it now